# Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations

### v1

@@ Discussion
Direct balancing weights in part tend to correspond to estimators with reduced variance because they focus only on balancing linear combinations of the set of basis functions $b(\bm{x}).$ However, if the CATE is not linear in the basis functions, the direct balancing weights estimator is generally biased for the ATE. This has motivated distributional balancing weights are derived to minimize the distance between each of the weighted treated group covariate distributions and the sample covariate distribution (Huling and Mak,, 2024). For distances that are integral probability metrics, the corresponding distributional balancing weights balance all functions in a given class (e.g, weights proposed by Huling and Mak, (2024) balance all functions in a particular Sobolev space). Thus, these weights balance a wider class of functional forms, which is desirable as $\mu_{z}(\bm{X})$ is often unknown. Let $\mathcal{D}$ a distributional distance measure and let $F_{n,z,\bm{w}}(\bm{x})$ be a weighted covariate distribution in treatment group $z$. Then distributional weights commonly take the follow form,
$\bm{w}_{n},=\text{argmin}_{\bm{w}}$ $\>\>\>\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}+\mathcal{D}\{F_{n,1,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}$
where $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ can also be added to minimization objective. While there is always a solution to this optimization problem, when there is a lack of overlap, the corresponding estimator may have inflated variance. To address this, Chen et al., (2024) propose weighting $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}+\mathcal{D}\{F_{n,1,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}$ by a hyperparameter $0\geq\alpha\geq 1$ and $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ by $1-\alpha$; increasing the relative weight of $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ tends to decrease variance but increase bias due to deviation from the target population. When $\alpha=0$, the corresponding weights target the ATO. However, this results in the entire covariate population changing for any $\alpha>0$. Thus, we propose the following distributional weights,
$\bm{w}_{n},=\text{argmin}_{\bm{w}}$ $\>\>\>\mathcal{D}[F_{n,0,\bm{w}}\{c(\bm{x})\},F_{n}\{c(\bm{x})\}]+\mathcal{D}[F_{n,1,\bm{w}}\{c(\bm{x})\},F_{n}\{c(\bm{x})\}]$
@@
where only a subset $c(\bm{x})$ of the covariate distribition is balanced to the sample distribution, but the entire covariate distribution is balanced between treatment and control. In addition, $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ could be replaced by $\mathcal{D}[F_{n,0,\bm{w}}\{g(\bm{x})\},F_{n,1,\bm{w}}\{g(\bm{x})\}]$ if desired. Alternatively, Mak and Joseph, (2018) define distributional distances based on a kernel that allows each covariate to impact the distance in different ways; this could be used within a standard distributional balancing weights set-up to weaken the impact of the $g(\bm{x})$ on the distributional distance between the weighted treatment groups and the sample population.
Additional guidance and implementation details for method
Intuition and guidance on selecting the covariates set only balanced between treated and control groups
@@
Selecting the $g(\bm{x})$ set requires researchers to evaluate the trade-offs between estimator variance and bias due to a modified $g(\bm{x})$ target population for their particular application. While these trade-offs are unavoidable when there are positivity violations, it is still desirable to select the $c(\bm{x})$ and $g(\bm{x})$ sets such that the corresponding estimator has minimal error with respect to an estimand of scientific interest, whether this the ATE or the average treatment effect for a modified population. Both subject matter knowledge as well as data driven techniques can be used to identify the $c(\bm{x})$ and $g(\bm{x})$ set; here, we provide further guidance on how a variety of factors and preferences may influence the selection of these two sets.
@@
As mentioned in the previous section, the corresponding estimator for Problem (5) has minimal error with respect to the ATE when $c(\bm{x})$ contains all treatment effect modifiers effect modifiers. When the ATE is of scientific interest, it is therefore desirable to determine the set $g(\bm{x})$ that 1) yields a solution to Problem (5); and 2) minimally modifies the treatment effect. There are a variety of scenarios where set of effect modifiers is likely to be smaller than the set of confounders. Of particular interest are scenarios with high dimensional confounders, which tend to yield overlap issues (D’Amour et al.,, 2021), but where the set of effect modifiers potentially have moderate to low dimension. Subject matter knowledge may indicate which covariates are unlikely to be effect modifiers this which covariates should be in the $g(\bm{x})$. A data-driven procedure could also be used to select $g(\bm{x})$ by estimating how much each covariate influences treatment effect heterogeneity; in fact, we propose an algorithm and corresponding model-based style estimator to do this Section 6.
@@
In other scenarios, either 1) all covariates of interest may be effect modifiers or 2) the size of $g(\bm{x})$ required to ensure a solution to Problem (5) may be large such that some effect modifiers need to be included in the set. From a design-based perspective, $g(\bm{x})$ is the set of covariates that may differ from the original sample population; thus in these cases, it is desirable to choose $g(\bm{x})$ such that the resulting weighted population is still meaningful within the given scientific context. In these cases, $g(\bm{x})$ could be chosen as the set of covariates for which their population is of less interest or relevance to the original research question. However, when there are high dimensional covariates in may be challenging to determine which subset yields a solution to Problem (5) and whose corresponding population may be modified without substantially altering the original research question. As described in Section 3.1, the set $g(\bm{x})$ needs to be predictive of treatment assignment $Z$ in order to substantially relax the positivity assumption. As such, there is some level of predictive capacity, $\eta$, that the covariate functions in $g(\bm{x})$ need to jointly satisfy in order for Problem (5) to yield a solution such that many possible $g(\bm{x})$ sets to select. However, selecting the covariate functions that are individually the most strongly predictive of treatment will tend to result in the smallest of such sets that have predictive capacity $\eta$. While selecting smallest $g(\bm{x})$ set that yields a solution to Problem (5) may not minimize the design-based components of estimator bias, it tends to both 1) simplify the interpretation of resulting the modified population; and 2) be more computationally efficient to operationalize. Thus, in Section 6 we proposed a design-based algorithm for identifying this set.
@@
Furthermore, both model- and design-based perspectives can be used together to select a set $g(\bm{x})$ that best matches the original research objective (e.g., $g(\bm{x})$ contains a mix of covariates that are either unlikely to be effect modifiers or highly predictive of $Z$). Therefore, while a lack of overlap forces researchers to confront trade-offs between estimator variance, statistical bias, and bias due population modification, we propose a variety of intuitive arguments and tools such that researchers can incorporate their subject-matter knowledge and research priorities when navigating these tradeoffs within our proposed optimization procedure.
@@
In Section 6 we proposed a model-based estimator. To preserve downstream inference we use the following cross-fit style estimator,
@@
$TEM_{j}=\left|\frac{1}{\sum_{i}X_{ij}^{2}}\sum_{i=1}^{n}X_{ij}\left\{(Y_{i}-\hat{\mu}_{Z_{i}}(\bm{X}_{i}))\frac{2Z_{i}-1}{Z_{i}\hat{e}(\bm{X}_{i})+(1-Z_{i})(1-\hat{e}(\bm{X}_{i}))}+\hat{\mu}_{1}(\bm{X}_{i})-\hat{\mu}_{0}(\bm{X}_{i})\right\}\right|,$
proposed by Boileau et al., (2025), as this treatment effect modification measure resulted in estimators with the lowest bias across a majority of the simulation scenarios, compared to the Hines et al., (2023) metric and a linear CATE model (supplementary Figure 2). In this estimator, $\hat{e}(\bm{x}),\hat{\mu}_{z}(\bm{x})$ are propensity score and outcome model estimates. Since propensity score estimates are likely to be extreme in scenarios with a lack of overlap we also explored $TEM^{*}_{j}=\text{abs}[\frac{1}{\sum_{i}X_{ij}^{2}}\sum_{i=1}^{n}X_{ij}\left\{\hat{\mu}_{1}(\bm{X}_{i})-\hat{\mu}_{0}(\bm{X_{i}})\right\}]$ which only requires outcome model estimates. In simulations, estimators corresponding to $TEM_{j}$ tended to perform similarly or better to estimators corresponding to $TEM^{*}_{j}$ (supplementary Figure 3). We use $TEM_{j}$ as the metric in Algorithm 1 as estimators corresponding to $TEM_{j}$ tended to perform similarly or better to estimators corresponding to $TEM^{*}_{j}$ (supplementary Figure 3).
@@
For the data generation propensity score model, $\alpha_{j}=0\pm a$ for $a=0.5/\gamma,1.33/\gamma,2.16/\gamma,3/\gamma$ for $j>0$. Here, $a$ determines the strength of relationship between a given covariate and treatment assignment where there are an equal number of coefficients generated with each of $a=0.5/\gamma,1.33/\gamma,2.16/\gamma,3/\gamma$. As $\gamma$ increases, propensity scores become more extreme and overlap decreases. We set $\mu_{0}(\bm{X})=0.25X_{1}+\cdots+0.25X_{p}$ and $\mu_{1}(\bm{X})=\sum_{i=0}^{p}\beta_{i}X_{i}$. We consider $\theta=0.25,0.75$ as the proportion of covariates are effect modifiers. We set $\beta_{0}=1$ and $\beta_{j}=0.25\pm c$ for $c=0.75\delta,0.6\delta,0.45\delta,0.3\delta,0$. Here, $c$ determines the strength of relationship between a given covariate and $\tau(\bm{X})$; for $(1-\theta)p$ coefficients $\beta_{j}=0$ and there are an approximately equal number of remaining coefficients set as each of $c=0.75\delta,0.6\delta,0.45\delta,0.3\delta$. Then, as $\delta$ increases, treatment effect heterogeneity increases or a given $\theta$. We generated 20 different potential types of covariates for each such scenario with all combinations of $\alpha_{j}$ and $\beta_{j}$ values. See supplementary Table 1 below for the $\gamma,\delta$ values used for each scenario.
@@
Table 1: Data generation overlap and treatment heterogeneity hyperparmeters for each simulation scenario.
Figure 1: Propensity score distribution for the simulation scenarios in Section 7. The scenarios are labeled on the top and right of the panels.
Figure 2: Comparison of estimator MSE with respect to the ATE for our proposed design-based estimator and model-based estimator (with the Boileau et al., (2025), Hines et al., (2023), and linear CATE model coefficient metrics) for both supplementary Algorithm 1 (not-adaptive) and Algorithm 1 (adaptive). The scenarios are labeled on the top and right of the panels.
Figure 3: Comparison of estimator MSE for our model-based estimator with the Boileau et al., (2025) and Hines et al., (2023) metrics calculated with and without estimated propensity scores. The scenarios are labeled on the top and right of the panels.
Figure 4: Estimator mean squared error (MSE) for all 40% treated simulation scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 5: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 20% treated and $p=20$. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 6: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 20% treated and $p=100$. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 7: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 40% treated and $p=20$. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 8: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 40% treated and $p=100$. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
@@
1:Inputs: Set of functions of covariates ordered according to some metric, $m$, $b_{(1)}(\bm{X}),\ldots,b_{(J)}(\bm{X})$
@@
3:while $w^{*}$ given $g(\bm{X})=b(\bm{X})^{-}$ does not exist do
@@
6:  Solve for $w^{(}$ in Problem (5) for $g(\bm{X})=b(\bm{X})^{-}$ and $c(\bm{X})=\{b_{(1)}(\bm{X}),\ldots,b_{(J)}(\bm{X})\}\setminus b(\bm{X})^{-}$
@@
9:   $b(\bm{X})^{-}$ $\triangleright$ Smallest set (given metric $m$) that yields a solution to Problem (5) when $g(\bm{X})=b(\bm{X})^{-}$
10:   $\{w^{*}\}_{i=1}^{n}$ $\triangleright$ Resulting weights from (5) with $g(\bm{X})=b(\bm{X})^{-}$
@@
Figure 9: Distribution of estimated propensity scores for the study on indwelling arterial catheters (IAC) with the data from the MIMIC-III database (Section 8)
@@
Figure 10: Treated and control group standardized mean differences (SMD) for all covariates for the study on indwelling arterial catheters (IAC) with the data from the MIMIC-III database (Section 8).
@@
Figure 11: Treated and sample population standardized mean differences (SMD) for all covariates for the study on indwelling arterial catheters (IAC) with the data from the MIMIC-III database (Section 8).
@@
Figure 12: Control and sample population standardized mean differences (SMD) for all covariates for the study on indwelling arterial catheters (IAC) with the data from the MIMIC-III database (Section 8).
@@
Figure 13: Distribution of estimated probabilities of being in the target population for transporting the health care hotspotting effect to a Midwestern academic health center population (Section 9)
@@
Figure 14: Treated and control group standardized mean differences (SMD) for all covariates for transporting the health care hotspotting effect to a Midwestern academic health center population (Section 9).
@@
Figure 15: Treated and target population standardized mean differences (SMD) for all covariates for transporting the health care hotspotting effect to a Midwestern academic health center population (Section 9).
@@
Figure 16: Control and target population standardized mean differences (SMD) for all covariates for transporting the health care hotspotting effect to a Midwestern academic health center population (Section 9).

### v2

@@ Discussion
Direct balancing weights tend to correspond to estimators with reduced variance in part because they focus only on balancing linear combinations of the set of basis functions $b(\bm{x}).$ However, if the CATE is not only linear in this set of basis functions, the direct balancing weights estimator is generally biased for the ATE. This has motivated the development of distributional balancing weights which are derived to minimize the distance between each of the weighted treatment group empirical distributions and the population empirical distribution (Huling and Mak, 2024). For distances that are integral probability metrics, the corresponding distributional balancing weights balance all functions in a given class (e.g, weights proposed by Huling and Mak (2024) balance all functions in a particular Sobolev space). Thus, these weights balance a much wider class of functions than direct balancing weights, which is desirable as $\mu_{z}(\bm{X})$ is often unknown. Let $\mathcal{D}$ a distributional distance measure, $F_{n}(\bm{x})$ be the empirical covariate distribution, and $F_{n,z,\bm{w}}(\bm{x})$ be the weighted empirical covariate distribution in treatment group $z$. Then distributional weights commonly take the follow form,
$\bm{w}_{n}=\text{argmin}_{\bm{w}}$ $\>\>\>\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}+\mathcal{D}\{F_{n,1,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}$
where $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ can also be added to minimization objective. While there is always a solution to this optimization problem, when there is a lack of overlap, the corresponding estimator may have inflated variance. To address this, Chen et al. (2024) propose weighting $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}+\mathcal{D}\{F_{n,1,\bm{w}}(\bm{x}),F_{n}(\bm{x})\}$ by a hyperparameter $0\geq\alpha\geq 1$ and $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ by $1-\alpha$; increasing the relative weight of $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ tends to decrease variance but increase bias due to deviation from the target population. When $\alpha=0$, the corresponding weights target the ATO. However, this results in the entire covariate population changing for any $\alpha>0$. Thus, we propose the following distributional weights,
$\bm{w}_{n}=\text{argmin}_{\bm{w}}$ $\>\>\>\mathcal{D}[F_{n,0,\bm{w}}\{c(\bm{x})\},F_{n}\{c(\bm{x})\}]+\mathcal{D}[F_{n,1,\bm{w}}\{c(\bm{x})\},F_{n}\{c(\bm{x})\}]$
@@
where the distribution of only a subset of the covariates (i.e., $c(\bm{x})$) is balanced to the empirical distribution, but all covariates are balanced between treated and control groups. In addition, $\mathcal{D}\{F_{n,0,\bm{w}}(\bm{x}),F_{n,1,\bm{w}}(\bm{x})\}$ could be replaced by $\mathcal{D}[F_{n,0,\bm{w}}\{g(\bm{x})\},F_{n,1,\bm{w}}\{g(\bm{x})\}]$ if desired. Alternatively, Mak and Joseph (2018) define distributional distances with a kernel that allows for the variable influence of each covariate on the distance. This could be used within a distributional balancing weights set-up to weaken the impact of the $g(\bm{x})$ on the distributional distance between the weighted treatment groups and the sample population.
Simulation study on estimator variance
We generate data using the same methods described in Section 6 and supplementary Section 6 with $\theta=100\%$ of covariates being treatment effect modifiers. Thus, we generate a total of 36 simulation scenarios that are all combinations p = 20, 100, 20% and 40% treated, three levels of overlap, and three levels of treatment effect heterogeneity. We compute a Monte Carlo estimate of the asymptotic variance of the direct balancing weights estimator (Corollary 4.1) and the asymptotic variance of our proposed estimator when $c(\bm{x})$ includes 75% and 25% of all covariates (Corollary 4.2). We also compute the empirical variance of these estimators when computing the weights with and without the additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$.
Figure 1: Estimated asymptotic variance estimates for all simulation scenarios. In the legend, “X% c(X)” indicates that X% of covariates are included in the $c(x)$ set for the PRTBW weighted estimators. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Across a majority of scenarios, our proposed estimators have smaller estimated asymptotic variance than the direct balancing weights estimator (supplementary Figure 1). However, when there is high treatment effect heterogeneity, our proposed estimator with only 25% of covariates in the $c(\bm{x})$ set tends to have larger estimated asymptotic variance than the direct balancing weights estimator, likely due to the large variance of $\tau_{g,\tilde{w}}(\bm{X},Z)$ compared to $\tau(\bm{X})$. Despite this, our proposed estimator with 75% of covariates in the $c(\bm{x})$ set has smaller asymptotic variance than the direct balancing weights estimator across almost all scenarios. Thus, while our proposed estimator will not always have smaller asymptotic variance than the direct balancing estimator, our proposed estimator does have similar or lower variance across many scenarios. The empirical estimator variance is similar the estimated asymptotic variance for all estimators across all scenarios which further validates the estimated asymptotic variances (supplementary Figure 2).
Figure 2: Estimated asymptotic variance estimates and empirical estimator variance for all simulation scenarios. In the legend, “X% c(X)” indicates that X% of covariates are included in the $c(x)$ set for the PRTBW weighted estimators. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
While the asymptotic results do not directly apply with the additional constraints of $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ to Problem (5), we can explore the behavior of the estimators in this case with empirical estimator variance. In many scenarios, there is no solution to the direct balancing optimization problem with these additional constraints such that empirical estimator variance cannot be computed (supplementary Figure 3). Across all scenarios where direct balancing weights could be calculated, our proposed estimators have smaller or similar empirical variance to the direct balancing weights estimator; when overlap is poor our proposed estimators have substantially smaller empirical variance. Therefore, our proposed estimators tend to outperform the direct balancing weights estimator in terms of variance across almost all scenarios when the weights are constrained to not extrapolate.
Figure 3: Empirical estimator variance for all simulation scenarios when including additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ in Problem (5). In the legend, “X% c(X)” indicates that X% of covariates are included in the $c(x)$ set for the weighted PRTBW estimators. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Additional guidance and implementation details
Guidance on selecting the $g(\bm{x})$ set
@@
Selecting the $g(\bm{x})$ set requires researchers to evaluate the trade-offs between estimator variance and estimator bias due to the modified $g(\bm{x})$ population for their particular application. While these trade-offs are unavoidable when there are positivity violations, it is still desirable to select the $c(\bm{x})$ and $g(\bm{x})$ sets such that the corresponding estimator has minimal error with respect to an estimand of scientific interest, whether this the ATE or the average treatment effect for a modified population. Both subject matter knowledge as well as data driven techniques can be used to identify the $c(\bm{x})$ and $g(\bm{x})$ set; here, we provide further guidance on how a variety of factors and preferences may influence the selection of these two sets.
@@
As mentioned in the previous section, the corresponding estimator for Problem (5) has minimal error with respect to the ATE when $c(\bm{x})$ contains all treatment effect modifiers effect modifiers. When the ATE is of scientific interest, it is therefore desirable to determine the set $g(\bm{x})$ that 1) yields a solution to Problem (5); and 2) minimally modifies the treatment effect. There are a variety of scenarios where set of effect modifiers is likely to be smaller than the set of confounders. Of particular interest are scenarios with high dimensional confounders, which tend to yield overlap issues (D’Amour et al., 2021), but where the set of effect modifiers potentially has moderate to low dimension. Subject matter knowledge may indicate which covariates are unlikely to be effect modifiers and thus which covariates should be in the $g(\bm{x})$. A data-driven procedure could also be used to select $g(\bm{x})$ by estimating how much each covariate influences treatment effect heterogeneity; in fact, we propose an algorithm and corresponding model-based estimator to do this Section 5.
@@
In other scenarios, either 1) all covariates of interest may be effect modifiers or 2) the size of $g(\bm{x})$ required to ensure a solution to Problem (5) may be large such that some effect modifiers need to be included in the set. From a design-based perspective, $g(\bm{x})$ is the set of covariates that may differ from the original sample population; thus in these cases, it is desirable to choose $g(\bm{x})$ such that the resulting weighted population is still meaningful within the given scientific context. In these cases, $g(\bm{x})$ could be chosen as the set of covariates for which their population is of less interest or relevance to the original research question. However, when there are high dimensional covariates in may be challenging to determine which subset yields a solution to Problem (5) and whose corresponding population may be modified without substantially altering the original research question. As described in Section 3.1, the set $g(\bm{x})$ needs to be predictive of treatment assignment $Z$ in order to substantially relax the positivity assumption. As such, there is some level of predictive capacity, $\eta$, that the covariate functions in $g(\bm{x})$ need to jointly satisfy in order for Problem (5) to yield a solution; thus, there are many possible $g(\bm{x})$ sets from which to select. However, selecting the covariate functions that are individually the most predictive of treatment will tend to result in the smallest of such sets that have predictive capacity $\eta$. While selecting the smallest $g(\bm{x})$ set that yields a solution to Problem (5) may not minimize the design-based component of estimator error, it tends to both 1) simplify the interpretation of the resulting modified population; and 2) be more computationally efficient to operationalize. Thus, in Section 5, we proposed a design-based algorithm for identifying this set.
@@
Furthermore, both model- and design-based perspectives can be used together to select a $g(\bm{x})$ set that best matches the original research objective (e.g., $g(\bm{x})$ contains a mix of covariates that are either unlikely to be effect modifiers or highly predictive of $Z$). Therefore, while a lack of overlap forces researchers to confront trade-offs between estimator variance, statistical bias, and bias due population modification, we propose a variety of intuitive arguments and tools such that researchers can incorporate their subject-matter knowledge and research priorities when navigating these tradeoffs within our proposed optimization procedure.
@@
In Section 5, we propose a model-based estimator. To preserve downstream inference we use the following cross-fit style estimator,
@@
[…]
proposed by Boileau et al. (2025), as this treatment effect modification measure resulted in estimators with the lowest bias across a majority of the simulation scenarios, compared to the Hines et al. (2023) metric and the coefficients from a linear CATE model (supplementary Figure 4).
Figure 4: Comparison of estimator MSE with respect to the ATE for our proposed design-based estimator and model-based estimator (with the Boileau et al. (2025), Hines et al. (2023), and linear CATE model coefficient metrics) for both supplementary Algorithm 1 (not-adaptive) and Algorithm 1 (adaptive). The scenarios are labeled on the top and right of the panels.
In this estimator, $\hat{e}(\bm{x}),\hat{\mu}_{z}(\bm{x})$ are propensity score and outcome model estimates. Since propensity score estimates are likely to be extreme in scenarios with a lack of overlap we also explored $TEM^{*}_{j}=\text{abs}[\frac{1}{\sum_{i}X_{ij}^{2}}\sum_{i=1}^{n}X_{ij}\left\{\hat{\mu}_{1}(\bm{X}_{i})-\hat{\mu}_{0}(\bm{X_{i}})\right\}]$ which only requires outcome model estimates. In simulations, estimators corresponding to $TEM_{j}$ tended to perform similarly or better to estimators corresponding to $TEM^{*}_{j}$ (supplementary Figure 5) such that we recommend and implement $TEM_{j}$ in applications of our methods.
Figure 5: Comparison of estimator MSE for our model-based estimator with the Boileau et al. (2025) and Hines et al. (2023) metrics calculated with and without estimated propensity scores. The scenarios are labeled on the top and right of the panels.
@@
For the data generation propensity score model, $\alpha_{j}=0\pm a$ for $a=0.5/\gamma,1.33/\gamma,2.16/\gamma,3/\gamma$ for $j>0$. Here, $a$ determines the strength of relationship between a given covariate and treatment assignment where there are an equal number of coefficients generated with each of $a=0.5/\gamma,1.33/\gamma,2.16/\gamma,3/\gamma$. As $\gamma$ increases, propensity scores become more extreme and overlap decreases. We set $\mu_{0}(\bm{X})=0.25X_{1}+\cdots+0.25X_{p}$ and $\mu_{1}(\bm{X})=\sum_{i=0}^{p}\beta_{i}X_{i}$. We consider $\theta=0.25,0.75$ as the proportion of covariates are effect modifiers. We set $\beta_{0}=1$ and $\beta_{j}=0.25\pm c$ for $c=0.75\delta,0.6\delta,0.45\delta,0.3\delta,0$. Here, $c$ determines the strength of relationship between a given covariate and $\tau(\bm{X})$; for $(1-\theta)p$ coefficients $\beta_{j}=0$ and there are an approximately equal number of remaining coefficients set as each of $c=0.75\delta,0.6\delta,0.45\delta,0.3\delta$. Then, as $\delta$ increases, treatment effect heterogeneity increases for a given $\theta$. Therefore, we generate 20 different potential types of covariates for each scenario (i.e., all combinations of $\alpha_{j}$ and $\beta_{j}$ values). See supplementary Table 1 for the $\gamma,\delta$ values used for each scenario.
Table 1: Overlap and treatment heterogeneity hyperparmeters used for data generation for each simulation scenario as described in Section 6 and supplementary Section 6.
@@
1:Inputs: The matrix $\mathbf{X}_{n\times p}$ and set of covariate functions ordered according to some metric, $m$, $b_{(1)}(\bm{X}),\ldots,b_{(J)}(\bm{X})$
@@
3:while $\{w^{*}_{i}\}_{i=1}^{n}$ given $g(\bm{X})=b(\bm{X})^{-}$ does not exist do
@@
6:  Solve Problem (5) for $g(\bm{X})=b(\bm{X})^{-}$ and $c(\bm{X})=\{b_{(1)}(\bm{X}),\ldots,b_{(J)}(\bm{X})\}\setminus b(\bm{X})^{-}$
@@
9:   $b(\bm{X})^{-}$ $\triangleright$ Smallest set that yields a solution to Problem (5) when $g(\bm{X})=b(\bm{X})^{-}$
10:   $\{w^{*}_{i}\}_{i=1}^{n}$ $\triangleright$ Resulting weights from (5) with $g(\bm{X})=b(\bm{X})^{-}$
Figure 6: Propensity score distribution for the simulation scenarios described in Section 6. The scenarios are labeled on the top and right of the panels.
Figure 7: Estimator mean squared error (MSE) for all 40% treated simulation scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 8: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 20% treated and $p=20$ scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 9: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 20% treated and $p=100$ scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 10: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 40% treated and $p=20$ scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
Figure 11: Estimator absolute percent bias with respect to the ATE and empirical standard deviation (SD) for 40% treated and $p=100$ scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing weights optimization problem did not exist for all datasets. The horizontal axis indicates levels of treatment effect heterogeneity while the simulation scenarios are labeled on the top and right of the panels.
@@
Figure 12: Distribution of estimated propensity scores for the study on indwelling arterial catheters (IAC) with data from the MIMIC-III database (Section 7.1)
@@
Figure 13: Treated and control group standardized mean differences (SMD) of all covariates for the study on indwelling arterial catheters (IAC) with data from the MIMIC-III database (Section 7.1).
@@
Figure 14: Treated and sample population standardized mean differences (SMD) of all covariates for the study on indwelling arterial catheters (IAC) with data from the MIMIC-III database (Section 7.1).
@@
Figure 15: Control and sample population standardized mean differences (SMD) of all covariates for the study on indwelling arterial catheters (IAC) with data from the MIMIC-III database (Section 7.1).
@@
Figure 16: Distribution of the estimated probability of being in the target population when transporting the “health care hotspotting” RCT effect to a Midwestern academic health center population (Section 7.2)
@@
Figure 17: Treated and control group standardized mean differences (SMD) of all covariates when transporting the “health care hotspotting” RCT effect to a Midwestern academic health center population (Section 7.2).
@@
Figure 18: Treated and target population standardized mean differences (SMD) of all covariates when transporting the “health care hotspotting” RCT effect to a Midwestern academic health center population (Section 7.2).
@@
Figure 19: Control and target population standardized mean differences (SMD) of all covariates when transporting the “health care hotspotting” RCT effect to a Midwestern academic health center population (Section 7.2).